OpenStudy (anonymous):

Find the standard deviation. Round to one more place than the data. 6, 5, 9, 13, 18, 7, 5, 20, 11, 15 A. 5.4 B. 1.2 C. 4.9 D. 5.1

OpenStudy (anonymous):

First find the mean (Sum up and divide by the number of data points) Then, for each data point, find the difference from the mean and square that. Add all of those up, divide by the number of data points, and take the square root! Ask me any questions that come up. =)

OpenStudy (anonymous):

i think im doing it all wrong!

OpenStudy (anonymous):

the first part im doing right its just the last part i need help on

OpenStudy (anonymous):

Okay, what did you get for the mean?

OpenStudy (anonymous):

after i added it all up i got 109 then i divided that by 10 so i got 10.9?

OpenStudy (anonymous):

Good =) Okay, now this part is a little bit tedious, but take each data point and do (10.9-x)^2 So, for the first one, it would be (10.9-6)^2 = (4.9)^2 = 24.01 Do that for each data point and add up all of the results.

OpenStudy (anonymous):

okay i got 267.79..

OpenStudy (anonymous):

so i did 24.01+ 34.81 + 3.61 + 4.41 + 50.41 + 15.21 + 34.81 + 82.81 + 0.9+ 16.81= 267.79 did i do it right?

OpenStudy (anonymous):

Hm... you're so close. I got 266.9. I'm not sure if you did one of those wrong of if it's just from rounding.

OpenStudy (anonymous):

i probably didn't round it off right! okay so after i get that result i square it?

OpenStudy (anonymous):

Last step is easy =) Divide by the number of data points. Then take the square root.

OpenStudy (anonymous):

5.4?

OpenStudy (anonymous):

Yes =)

OpenStudy (anonymous):

ahhh thank you so much :D

OpenStudy (anonymous):

My pleasure =D